### v1

# Coherent control through phonon anharmonicity

Abstract
Anharmonic lattice vibrations play a key role in many physical phenomena. They govern the heat conductivity of solids, strongly affect the phonon spectra, play a prominent role in soft mode phase transitions, allow ultrafast engineering of material properties and more. The most direct evidence for anharmonicity is to measure the oscillation frequency changing as a function of the oscillation amplitude. For lattice vibrations, this is not a trivial task, and anharmonicity is probed indirectly through its effects on thermodynamic properties and spectral features or through coherent decay of one mode to another. However, measurement of the anharmonicity of a single Raman mode is still lacking. We show that ultrafast double pump-probe spectroscopy could be used to directly observe frequency shifts of Raman phonons as a function of the oscillation amplitude and disentangle the coherent contributions from quasi-harmonic sources such as temperature and changes to the carrier density in the thermoelectric materials SnTe and SnSe. Our results have dramatic implications for the material engineering of future thermoelectrics. Moreover, our methodology could be used to isolate the basic mechanisms driving optically induced phase transitions and other nonlinear phenomena based on their unique timestamps.
Raymond and Beverly Sackler School of Physics and Astronomy, Tel-Aviv 69978, Israel
The properties of phonons and their interplay with other constituents of a solid play a crucial role in determining material properties. One such property is the phonon anharmonicity, which determines the lattice thermal conductivity, and understanding it is critical for the development and engineering of new thermoelectric and energy materials[1, 2]. Phonon anharmonicity could be indirectly probed by measuring the change in scattering or thermodynamic properties as a function of external stimuli such as temperature and pressure. For example, it can be observed as line-width broadening and changes to the phonon frequency, as well as variations in thermal conductivity or expansion[3, 4] as the temperature changes. Alternatively, its effects on the phonon spectrum could be probed through an inelastic scattering of neutrons[1] and X-rays[5]. Ultrafast pump-probe spectroscopy has also been used to observe anharmonic coupling between phonon modes. For example, Teitelbaum et al.[6] showed that the phonon in excited bismuth decays to pairs of acoustic modes with opposite momenta which drove temporal oscillations at the corresponding frequency. Similarly, off-diagonal peaks in two-dimensional spectroscopy measure the coupling between different vibrational modes of molecules and solids[7, 8, 9, 10]. Another example is the nonlinear coupling between infrared-active modes and symmetric Raman modes was observed by Först et al.[11], pioneering the field of nonlinear phononics. The common feature of all of these experiments is that they probed anharmonic coupling between different vibrational modes. However, to our knowledge, no experiment has yet measured the frequency shifts of a single mode resulting from variations in the oscillation amplitude.
Anharmonicity is particularly important in thermoelectric materials and controlling it, as well as the frequencies, lifetime, and dispersion of both acoustic and optical modes are utilized as control knobs for thermoelectric efficiency.[12] We chose to work with SnTe, a promising thermoelectric material[13], whose anharmonicity is currently a topic of extensive study[14, 2, 15, 16]. Additionally, SnTe has recently gained further attention thanks to its topological crystalline insulator properties[17]. In this work, we directly measure the anharmonicity of a single phonon mode by utilizing double pump-probe spectroscopy, a technique used in the past to control phase transitions using light[18, 19, 20], to coherently control phonon oscillations[6, 21, 22], and more. Our findings reveal significant frequency shifts resulting from variations in oscillation amplitude. This method allows us to disentangle the contributions of different physical effects to ultrafast measurements using their temporal fingerprint such as thermal, electronic, and anharmonic effects on phonon behavior.
To excite coherent oscillations of the mode, we utilize the ”displacive excitation of coherent phonons” (DECP) process[23]. Ultrashort laser pulses (70 and 126 fs) at 750 nm, resonant with a transition between Te and Sn bands, are incident on the sample, resulting in photo-excitation of electrons from the Te anion to the Sn cation. This reduces the charge difference between the ions, resulting in a softening of the chemical bond between them. The charge transfer shifts the equilibrium position of the ions closer to one another causing them to oscillate around the new equilibrium positions.
Figure 1(a) shows the relative change of sample reflectivity as a function of the time delay from a single pump excitation with an incident fluence of 0.68 (which translates to excitation density of [24]). The signal could be decomposed to a decaying bi-exponential background overlayed with a damped oscillatory component. The colored solid lines in Figure 1(a) show the decomposition of the signal to its different components. We note that the oscillatory component displays a cosine-like behavior in agreement with the DECP mechanism[23]. The sign of the oscillatory component near time zero is positive. Understanding that the DECP mechanism in SnTe is initiated by bond softening enables us to use the positive sign of at time zero as a calibration, linking bond softening and the positive sign of the oscillatory component of . The fast exponential component can be associated with electrons decaying back to their ground state due to the similar timescales observed using time-resolved photoemission[25]. The slower exponential component can be associated with thermal relaxation effects between the electronic and phononic baths. Estimates of the expected time constants from known thermal conductivity and heat capacity of SnTe are shown in supplementary section 4. Figure 1(b) shows the fluence dependence of the signal of SnTe. The most striking aspect of the data is the redshift the oscillations exhibit towards higher fluences indicating strong softening of the mode as the fluence is increased. Dashed lines are used as a guide to the eye to mark the shift in the oscillation peaks as the fluence is varied. The shifts in peak positions are very small for the zero and first peaks and increase for the later peaks which is a signature that the shifts originate from changes in the oscillation frequency rather than its phase. Figure 1(c) shows the frequency that was extracted by fitting the data in Figure 1(b) as a function of the applied fluence. The magnitude of the phonon softening at the highest fluence is almost 20% lower than its value in equilibrium. An extrapolation of the oscillation frequency to zero fluence gives a frequency of corresponding to 122 cm-1, compatible with the known Ag phonon of SnTe[26] and our Raman measurements (inset), marked by a yellow star.
Optical phonon softening as a result of increasing laser fluence was observed in several compounds in the past[27, 28, 29, 30, 31]. The most relevant example for our work is SnSe where the phonon softening was associated with softening of the bond due to the photo-excitation of electrons[27]. Similar phenomena were also observed in bismuth where the nature of the phonon softening, whether electronic or anharmonic, was debated in a series of papers[28, 29, 30]. From these examples alone, it is clear that phonon softening as a function of the laser fluence can originate from various sources. This is due to the multitude of effects that occur when the pump pulse interacts with the sample, including heating, photo-excitation of electrons, and the excitation of coherent phonons. Ideally, one would isolate the contributions of the different effects by tuning them individually. While this is possible for infrared active modes via resonant excitation[32, 33, 34, 35] it is a very challenging task for the amplitude Raman active modes since their excitation is typically a non-resonant process such as DECP[23] which inherently involves an electronic excitation or a sudden temperature increase. Two unique cases are the experiments by Maehrlein et al[36] and Giorgianni et al[37] who were able to resonantly pump a Raman active phonons in diamond and respectively by using sum frequency of two low energy phonons in the mid-infrared and THz regimes respectively. In the case of diamond this was made possible thanks to the very high frequency (40 THz) of the particular mode and its unusually high Raman activity and for the case of it was made possible thanks to a very strong THz field and similarly strong Raman activity. A double pump experiment by Murry et al[30] shows that in bismuth fluence dependent softening originates from the coupling of the phonon to the electron-hole plasma by showing that the phonon frequency does not change as a function of the pump-pump delay. In our experiment, we go beyond what was previously demonstrated and show that the distinct transient characteristics of structural, electronic, and thermal effects serve as unique fingerprints to isolate their impacts on phonon anharmonicity using double pump-probe spectroscopy.
The incident fluences of the two pump beams were set to 2.822 for the leading pump and 0.571 for the trailing pump. These correspond to excitation densities of and respectively. Their respective single pump traces are shown in Figure 2(a). The observed frequencies are 3.14 THz and 3.48 THz, corresponding to the pulses that will act as leading and trailing pump pulses in the double pump-probe experiment. Our optical setup allows us to isolate the portion of originating from the trailing pump by modulating it with a mechanical chopper. The efficiency of the isolation process is demonstrated in Figure 2(b). The time delay between the two pump pulses () was set to 3.31 ps, and the arrival time of the probe pulse was scanned relative to the two pump pulses. The blue trace shows the measurement when the chopper modulates both pump pulses, whereas the orange trace shows the measurement where only the trailing pump was. The measurement was repeated for different values of and the results are shown in Figure 2(c). The oscillation frequency of changes with , as could be seen by the horizontal shift of the oscillation peaks. However, in contrast to the monotonically increasing mode softening, which was observed as a function of laser fluence (Figure 1), a nonmonotonic oscillatory behavior as a function of is observed.
To closely examine the origin of the nonmonotonic behavior, we plot in Figure 3 the frequency of the oscillations as a function of (detailed analysis in supplementary section 3). Figure 3(a) focuses on the long timescale features in the data. Significant softening is observed which follows a double exponential behavior as a function of with amplitudes and time constants of , , , . The long timescale of the second exponent is characteristic of thermal relaxation and we therefore attribute this contribution to quasi-harmonic effects[38] induced by the heating of the leading pump. The fast exponent is characteristic of electronic relaxation and is related to the relaxation of the photo-excited electrons back to their original states. This timescale is comparable to electronic relaxation timescales measured in the past by time-resolved photo-emission spectroscopy[25]. The softening characteristic of the electronic feature is expected since we know that photo-excitation acts to soften the chemical bonds between the Te anions and Sn cations. Moreover, it is compatible with the behavior calculated and observed in the sister compound SnSe[27]. Figure 3(b) focuses on short timescales of the and shows a distinct periodic non-sinusoidal behavior. The oscillatory component of the trace obtained when only the leading pump was incident on the sample is plotted for comparison. The periods and time constants of both signals match, indicating that the periodic component of the frequency shift is caused by displacing the atoms along the normal coordinate serving as a direct measurement for the modes anharmonicity. The magnitude of the oscillations reaches values as large as 0.3[THz] which is larger than both the thermal and the electronic effects. Moreover, this value is larger than what was observed for a similar value of the fluence in the single pump experiment. The sign of the periodic signal is correlated with that of the sinusoidal component of the signal. A positive and extension of the mode is accompanied by phonon softening while a negative is accompanied by hardening and contraction of the mode. This is in agreement with our previous understanding (Figure 1) that the softening of the mode is correlated with a positive sign of the oscillations at time zero where the DECP process acts to extend the mode. The periodic signal is asymmetric between its rising and descending parts. The rising side is slow, smooth, and seems to have an inflection point at the times where the signal crosses zero. In contrast, the descending part is sharp, and its exact features are beyond our experimental resolution which is limited by the duration of the two pump pulses. The asymmetry between the rising and descending parts of the data indicates that it is not only the position of the atoms along the phonon coordinate that determines the observed frequency but also the direction of their momentum at the time of photo-excitation. The origin of symmetry breaking is the directional nature of the DECP process, i.e., mode extension accompanied by softening. Similar asymmetric line-shapes were observed in the phase of coherent phonons in diamond[22] and in the phonon lifetime in bismuth[39] which were also excited using the DECP process. We note that while the similar line-shapes are characteristic of the DECP process, lattice anharmonicity was not observed in either bismuth or diamond. Similar measurements were taken using different fluences for the leading and trailing pump pulses. The full datasets and their analysis are provided in supplementary section 5 and show similar results. The signals vary in their amplitude and characteristic timescales which demonstrates the potential of the double pump method to control material properties.
SnTe has the rocksalt structure, therefore its fully symmetric Ag mode is similar to a uniform expansion of the crystal which also happens as a function of increased temperature. Therefore, the observed changes in frequency could be thought of as the coherent and thermal versions of the lattice expansion effect. To test whether the double pump method can identify anharmonicity along less symmetric phonon modes we performed similar measurements on SnSe, a low-symmetry variant of SnTe. Figure 4(a) shows the signal taken with single pump excitation. It is clear from the beat-like pattern, that more than a single frequency is present in the signal corresponding to more than one phonon in agreement with previous measurements of SnSe[27]. Figure 4(b) shows as a function of time delay and . Unlike in Figure 2(b), the presence of multiple frequencies in the data makes it challenging to decipher the anharmonic effects directly from the plot. To decrease the amount of fitting parameters we first subtract the exponential background and fit the data to a combination of damped oscillations to extract the mode frequencies as a function of . The results for the two most dominant phonons are shown in Figure 4 (c-d). Both modes were observed previously in single pump experiments[27] and are known to soften with increased fluence. However, as is increased both modes exhibit periodic modulations at the frequency excited by the leading pump for the same mode. In contrast to SnTe, the lower symmetry of SnSe differentiates between lattice expansion and motion along the crystals modes, and Huang et.al[27] even showed that their excitation is entwined with a novel lattice instability towards a higher symmetry phase.
Our experiment serves as a direct measurement of the anharmonicity of a single phonon mode originating from deviations from the parabolic shape of the potential for the ions. This is in contrast to line broadening and coherent coupling between different modes which are typically used to characterize anharmonicity. Our study demonstrates that double pump-probe spectroscopy is a powerful tool for disentangling the thermal, electronic, and anharmonic contributions to phonon frequency shifts. By applying this technique, we have been able to gain detailed insights into the complex interplay of these factors in SnTe and SnSe. Importantly, the applicability of our approach to disentangling electronic from thermal effects in ultrafast measurements is broader than the application discussed here. For example, light-induced phase transitions are tuned as a function of the pump fluence[40]. In many cases the nature of the transition, whether of thermal or electronic origins is debated. Typically, a theoretical understanding of the particular system at hand is required in addition to the experimental data to reach a definitive conclusion. We show that the double pump technique separates between thermal and electronic effects by their distinct temporal behavior and can greatly aid in solving these debates.
SnTe single crystals were grown using the vapor transport technique without the aid of a transport agent similar to the procedure described in Ref[41]. Stoichiometric amounts of Sn(99.995% Sigma Aldrich) and Te(99.999% Thermoscientific) were weighed and thoroughly mixed inside an alumina crucible. The starting materials were sealed under vacuum in a 20 cm long quartz ampule which was placed in a gradient tube furnace. The starting materials were positioned in the hot zone of the furnace. The source and growth zones of the furnace were ramped to 797°C and 747°C respectively at a rate of 25°C/hour. The synthesis took place for 1 week after which large single crystals were found in the cold zone of the tube. A similar procedure was used to grow SnSe. Selected crystals were ground to a fine powder using a mortar and pestle which was used to characterize the structure of the crystals using powder X-ray diffraction (see supplementary section 2).
Detailed information about the optical experiment, not provided in the main text, is shown in supplementary section 1.
A.R. acknowledges support from the Zuckerman Foundation, and the Israel Science Foundation (Grant No. 1017/20). This research was supported by the Ministry of Innovation, Science and Technology, Israel. Gili Scharf acknowledges support from the Israeli Clore fellowship.
The data that support the findings of this study are available from the corresponding author upon reasonable request.
All authors contributed to the production of this manuscript.
The authors declare no competing financial interests.
Correspondence and requests for materials should be addressed to Alon Ron (email: alonron@tauex.tau.ac.il).